OpenStudy (anonymous):

What is the max. value of T(x,y,z) = 1+xy+yz given that x,y and z all lie in a unit sphere?

OpenStudy (apoorvk):

pkay then, let the unit sphere be: x^2 + y^2 +z^2 =1 now you can maximize 1+xy+yz for the above constraint now. 

OpenStudy (anonymous):

Use spherical coordinates as in http://tutorial.math.lamar.edu/Classes/CalcII/SphericalCoords.aspx Replace x, y, z by their values. Since we are on the unit sphere \[ \rho =1\] The function to be maximized becomes \[ h(\theta,\phi)= \sin (\theta ) \cos (\theta ) \sin ^2(\phi )+\sin (\theta ) \sin (\phi ) \cos (\phi )+1 \] It is gradient is \[ gradienth(\theta,\phi)=\left\{-\sin ^2(\theta ) \sin ^2(\phi )+\cos ^2(\theta ) \sin ^2(\phi )+\cos (\theta ) \sin (\phi ) \cos (\phi ),-\sin (\theta ) \sin ^2(\phi )+\sin (\theta ) \cos ^2(\phi )+2 \sin (\theta ) \cos (\theta ) \sin (\phi ) \cos (\phi )\right\} \] Use your favorite symbolic system to find that the gradient is zero at 15 points. Four of them give a maximum of h \[ 1+\frac{1}{\sqrt{2}} \]

OpenStudy (anonymous):

The first point is \[ \left\{\frac{1}{2},\frac{1}{\sqrt{2}},\frac{1}{2}\right\} \] The second point is \[ \left\{-\frac{1}{2},-\frac{1}{\sqrt{2}},-\frac{1}{2}\right\} \] The last two points are the same of the first two.

OpenStudy (phi):

Here is how to do it using lagrange multipliers.